# Learning Fluid-Structure Interaction Dynamics with Physics-Informed Neural Networks and Immersed Boundary Methods

### v2

@@ 1 Introduction
Recent advancements in machine learning have introduced new methods for capturing complex fluid dynamics and structural responses with high computational efficiency [5, 6, 7, 8, 9, 10, 11, 12]. PINNs offer a compelling approach by integrating observational data with governing physical laws to solve partial differential equations (PDEs). PINNs have been successfully applied to nonlinear dynamics [13, 14, 15, 16, 17]. Several studies have explored PINNs for FSI problems, showing promising results in predicting flow fields and structural responses. Integrating PINNs with IBM has demonstrated the potential to combine the strengths of both approaches, enabling simple and efficient modeling without excessive computational overhead [18]. The IB-PINN frameworks [19, 20] apply penalty terms to enforce interface conditions but remain limited to stationary geometries and steady flows. More complex dynamics studies using operator learning and reduced-order models [7, 5] often lack explicit interface physics and require extensive training data. Moving boundary problems with region-specific loss partitioning and improved efficiency was proposed by [18], which still faces challenges with pressure recovery and sharp interface resolution.
@@ 3.1 Problem Setup
The computation domain, denoted as $\Omega$, is a two-dimensional square cavity defined by $\Omega$ $=[0,1]\times[0,1]$. The immersed structure is a disc with a radius of 0.2, initially centered at the point $\mathbf{x}=(0.6,0.5)$. We consider a uniform discretization with a grid $(N_{x},N_{y})=(100,100)$ and a time span $0\leq t\leq 10$, with a time step of 0.01. During this interval, the disc completes slightly more than one full rotation within the cavity. The boundary $\Gamma_{1}$ represents the top Dirichlet boundary condition of the cavity with velocity tangent to this side, while $\Gamma_{0}$ denotes the other three stationary sides. The Reynolds number of the flow is set to $Re=100$. During the simulation, the flow induced by the driven lid causes the structure to come nearly into contact with the moving upper boundary of the domain, as illustrated in Fig. 1.
@@
[…]
@@ 4.1 Single-FSI Network
$\mathcal{L}_{\text{bc}_{1}}(\theta)$ $=\mathcal{L}_{\text{bottom, right, left}}=\text{MSE}\left[\mathcal{L}(\hat{u})+\mathcal{L}(\hat{v})\right],$ (3f)
@@ 4.2 Eulerian-Lagrangian Network
$\mathcal{L}_{\text{bc}_{1}}(\theta_{1})$ $=\mathcal{L}_{\text{bottom, right, left}}=\text{MSE}\left[\mathcal{L}(\hat{u})+\mathcal{L}(\hat{v})\right],$ (4e)
@@ 4.3 Model Details
M3 Eulerian-Lagrangian (Fig. 3) Eulerian [3, 300, 300, 300, 3] 365105 Tanh
@@
The Tanh-activated configurations are implemented as an MLP. These models adopt a fully connected feedforward network, where each layer is initialized using the Xavier normal distribution to ensure stable gradients during training. The model normalizes inputs to the range $[-1,1]$, to enhance training stability when using tanh activations. In contrast, the learnable B-spline basis functions for activation are implemented using the Kolmogorov–Arnold Networks (KAN) [29], which integrate adaptive spline representations into the network’s functional form. Each layer consists of a combination of a standard linear transformation with a base activation (e.g., SiLU), and a spline-based nonlinearity defined as:
@@
where, $\{\xi_{i}\}$ are the knot points that define the spline, which are dynamically updated during training based on the distribution of input values.
@@ 4.4 Training Setup
To ensure fair evaluation of the models, each model was trained for 60,000 iterations, using a decay step of 1,000 and a decay rate of 0.99. We employed PyTorch’s Adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$. A Sobol sequence was used for the initial training dataset selection, and mini-batch gradient descent was performed with a batch size of 128, where samples were randomly selected from the Sobol-generated dataset to enhance model generalization. The training dataset included 0.005%, 0.05% of the fluid and solid interface data, respectively, ensuring efficient and representative sampling of boundary conditions and the solid interface. Furthermore, the training dataset and settings were standardized across all models to ensure a fair comparison and consistency in performance evaluation.
@@
Relative $\text{Relative }L_{2}\text{ norm}=\frac{\|y-\hat{y}\|_{2}}{\|y\|_{2}}*100\%$ (7)
@@ 5.4.2 Benefits of domain specialization
Our Eulerian-Lagrangian architecture addresses these limitations by decoupling the Eulerian and Lagrangian predictions into distinct specialized networks. This separation enables each network to focus exclusively on its respective domain-specific physics without interference from competing objectives. The fluid network can optimize for smooth, continuous flow fields governed by the Navier-Stokes equations, while the interface network can specialize in capturing the sharp gradients and boundary conditions characteristic of solid-fluid interfaces. This approach offers several key advantages over the Single-FSI architecture: (a) Each specialized network can accommodate the statistical distribution of its respective domain, reducing the impact of variable scaling issues during training. (a) Network capacity can be allocated more efficiently, with complexity distributed according to the specific modeling challenges of each domain. (c) The dedicated interface network can focus entirely on enforcing boundary conditions and capturing sharp transitions, leading to significantly more accurate velocity and pressure predictions at the critical fluid-solid boundary.

### v3

@@ 1 Introduction
Recent advancements in machine learning have introduced new methods for capturing complex fluid dynamics and structural responses with high computational efficiency [5, 6, 7, 8, 9, 10, 11, 12, 13]. PINNs offer a compelling approach by integrating observational data with governing physical laws to solve partial differential equations (PDEs). PINNs have been successfully applied to nonlinear dynamics [14, 15, 16, 17, 18]. Several studies have explored PINNs for FSI problems, showing promising results in predicting flow fields and structural responses. Integrating PINNs with IBM has demonstrated the potential to combine the strengths of both approaches, enabling simple and efficient modeling without excessive computational overhead [19]. The IB-PINN frameworks [20, 21] apply penalty terms to enforce interface conditions but remain limited to stationary geometries and steady flows. More complex dynamics studies using operator learning and reduced-order models [7, 5] often lack explicit interface physics and require extensive training data. Moving boundary problems with region-specific loss partitioning and improved efficiency was proposed by [19], which still faces challenges with pressure recovery and sharp interface resolution.
@@ 3.1 Problem Setup
The computation domain, denoted as $\Omega$, is a two-dimensional square cavity defined by $\Omega$ $0,1]\times[0,1]$. The immersed structure is a disc with a radius of 0.2, initially centered at the point $\mathbf{x}=(0.6,0.5)$. We consider a uniform discretization with a grid $(N_{x},N_{y})=(100,100)$ and a time span $0\leq t\leq 10$, with a time step of 0.01. During this interval, the disc completes slightly more than one full rotation within the cavity. The boundary $\Gamma_{1}$ represents the top Dirichlet boundary condition of the cavity with velocity tangent to this side, while $\Gamma_{0}$ denotes the other three stationary sides. The Reynolds number of the flow is set to $Re=100$. During the simulation, the flow induced by the driven lid causes the structure to come nearly into contact with the moving upper boundary of the domain, as illustrated in Fig. 1.
@@
$\frac{\partial\mathbf{d}_{s}(\xi(t,s))}{\partial t}$ $=\mathbf{u}_{s}(\xi(t,s))=\int_{\Omega}\mathbf{u}_{f}(t,\mathbf{x})\delta(\mathbf{x}-\xi(t,s))\,d\mathbf{x}=\mathbf{u}_{f}(\xi(t,s)),$ (1b)
@@ 4.1 Single-FSI Network
[…]
@@ 4.2 Eulerian-Lagrangian Network
[…]
@@ 4.3 Model Details
M3
Eulerian-Lagrangian
(Fig. 3)
Eulerian [3, 300, 300, 300, 3] 365105 Tanh
@@
The Tanh-activated configurations are implemented as an MLP. These models adopt a fully connected feedforward network, where each layer is initialized using the Xavier normal distribution to ensure stable gradients during training. The model normalizes inputs to the range [1,1][-1,1], to enhance training stability when using tanh activations. In contrast, the learnable B-spline basis functions for activation are implemented using the Kolmogorov–Arnold Networks (KAN) [30], which integrate adaptive spline representations into the network’s functional form. Each layer consists of a combination of a standard linear transformation with a base activation (e.g., SiLU), and a spline-based nonlinearity defined as:
@@
where $\{\xi_{i}\}$ are the knot points that define the spline, which is dynamically updated during training based on the distribution of input values.
@@ 4.4 Training Setup
To ensure a fair evaluation of the models, each model was trained for 60,000 iterations, using a decay step of 1,000 and a decay rate of 0.99. We employed PyTorch’s Adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$. A Sobol sequence was used for the initial training dataset selection, and mini-batch gradient descent was performed with a batch size of 128, where samples were randomly selected from the Sobol-generated dataset to enhance model generalization. The training dataset included 0.005%, 0.05% of the fluid and solid interface data, respectively, ensuring efficient and representative sampling of boundary conditions and the solid interface. Furthermore, the training dataset and settings were standardized across all models to ensure a fair comparison and consistency in performance evaluation.
@@
$\text{Relative }L_{2}\text{ norm}=\frac{\|y-\hat{y}\|_{2}}{\|y\|_{2}}*100\%$ (7)
@@ 5.4.2 Benefits of domain specialization
Our Eulerian-Lagrangian architecture addresses these limitations by decoupling the Eulerian and Lagrangian predictions into distinct specialized networks. This separation enables each network to focus exclusively on its respective domain-specific physics without interference from competing objectives. The fluid network can optimize for smooth, continuous flow fields governed by the Navier-Stokes equations, while the interface network can specialize in capturing the sharp gradients and boundary conditions characteristic of solid-fluid interfaces. This approach offers several key advantages over the Single-FSI architecture: (a) Each specialized network can accommodate the statistical distribution of its respective domain, reducing the impact of variable scaling issues during training. (b) Network capacity can be allocated more efficiently, with complexity distributed according to the specific modeling challenges of each domain. (c) The dedicated interface network can focus entirely on enforcing boundary conditions and capturing sharp transitions, leading to significantly more accurate velocity and pressure predictions at the critical fluid-solid boundary.